OpenStudy (anonymous):

Find the values of x in the following equation: |10x + -9| = 59

jhonyy9 (jhonyy9):

10x-9=-59 and 10x-9=59 10x=-50 10x=68 x= -50/10 x=68/10 x=-5 x=6,8